Find the value of z if 4 + z + 2 = 20 + 2z
Answer by marilynh(7)

4 + z + 2 = 20 + 2z
6 + z = 20 + 2z

You need to move the variable z to one side and the numbers to the other side of the = sign.

Subtract z from both sides:
6 = 20 + z

Subtract 20 from both sides:
-14 = z